# Disambiguate First Parse Later: Generating Interpretations for Ambiguity Resolution in Semantic Parsing

### v1

@@ Abstract
Handling ambiguity and underspecification is an important challenge in natural language interfaces, particularly for tasks like text-to-SQL semantic parsing. We propose a modular approach [^1] Code: github.com/saparina/disambiguate-then-parse that resolves ambiguity using natural language interpretations before mapping these to logical forms (e.g., SQL queries). Although LLMs excel at parsing unambiguous utterances, they show strong biases for ambiguous ones, typically predicting only preferred interpretations. We constructively exploit this bias to generate an initial set of preferred disambiguations and then apply a specialized infilling model to identify and generate missing interpretations. To train the infilling model, we introduce an annotation method that uses SQL execution to validate different meanings. Our approach improves interpretation coverage and generalizes across datasets with different annotation styles, database structures, and ambiguity types.
Disambiguate First Parse Later: Generating Interpretations for
@@
Irina Saparina  and Mirella Lapata Institute for Language, Cognition and Computation, School of Informatics, University of Edinburgh i.saparina@sms.ed.ac.uk   mlap@inf.ed.ac.uk
@@ 1 Introduction
While state-of-the-art large language models (LLMs) demonstrate remarkable performance on tasks like question answering, text-to-SQL parsing, and natural language inference, recent studies (Liu et al., 2023; Floratou et al., 2024) have shown they are not adept at handling ambiguity. They display systematic biases in their choice of interpretation (Kamath et al., 2024; Stengel-Eskin et al., 2024; Saparina and Lapata, 2024a), typically default to a single interpretation when multiple ones exist.
@@
Figure 2: Disambiguate first parse later: we disambiguate database questions into natural language interpretations (generate initial set and fill in missing ones), and then a parser translates each interpretation into SQL.
@@ 2.1 Problem Formulation
In our text-to-SQL task, grammar $G$ defines valid SQL queries for a given database schema $S$. The database schema and table descriptions provide context $\mathcal{C}$ that can help disambiguate some queries. We focus on questions that remain ambiguous even when $\mathcal{C}$ is known, i.e., there exist multiple valid SQL queries $\{e_{1},...,e_{n}\}$ that respect $S$ and produce different result sets $\llbracket e_{i}\rrbracket$ when executed on the database. In the example in Figure 1, the question $u$ = “return the rating of each hotel” remains ambiguous for a given schema $S$ (see top table), as “rating” could refer to star ratings and guest reviews even when database content is known.
@@ 2.2 Natural Language Interpretations for Explicit Disambiguation
We propose to resolve ambiguity prior to generating SQL expressions with natural language interpretations. More formally, for ambiguous utterance $u$, we first produce a set of unambiguous natural language interpretations $\{\hat{u}_{1},...,\hat{u}_{n}\}$, where each $\hat{u}_{i}$ captures one meaning of $u$. We then map each $\hat{u}_{i}$ to formal expression $e_{i}$. For example, given $u$ = “return the rating of each hotel”, our goal is to produce $\hat{u}_{1}$ = “How many stars were assigned to each hotel?”, $\hat{u}_{2}$ = “How did the customers review each hotel?” and $\hat{u}_{3}$ = “Show me the guest scores and star rating of each hotel” and mapped them to corresponding SQL queries $e_{1}$, $e_{2}$ and $e_{3}$ (see Figure 1). This approach has several advantages:
  • Unambiguous interpretations $\hat{u}_{i}$ are easier to map to formal expressions, as they can be handled by standard semantic parsers (e.g., existing text-to-SQL models);
@@ 3.2 Evaluation Metrics
Our primary metric is Full Interpretation Coverage, which measures the proportion of examples where all valid SQL interpretations are present in the output. This metric aligns with the “All Found” metric from Ambrosia and “BothInTopK” from AmbiQT. We also report Single Interpretation Coverage, i.e., the proportion of examples where at least one valid interpretation is found (“EitherInTopK” from AmbiQT). We compare SQL queries based on their execution results.
@@ 4 Experimental Results
Table 1 is split into two sections. The first one presents end-to-end approaches, which attempt to directly predict multiple SQL queries for ambiguous questions. We report results for zero-shot prompting, few-shot prompting, and end-to-end fine-tuning. For prompting, we follow Saparina and Lapata (2024a) and explicitly instruct the model to generate multiple SQL queries if the question is ambiguous. For few-shot prompting, we sample 3 random examples from the corresponding dataset. The second section in Table 1 lists methods which use natural language interpretations to disambiguate first and then rely on text-to-SQL parsing of unambiguous questions. We report results for generating all possible interpretations through LLM prompting (which corresponds to our method without infilling), applying self-correction to this approach (details in Appendix F), a fine-tuning method which is trained to generate reference interpretations instead of SQL queries, and our proposal which uses infilling to augment the set of default interpretations. We use an instruction-tuned Qwen2.5-Coder 32B for text-to-SQL generation.
AmbiQT Ambrosia
Method Single Full Single Full
End-to-End Text-to-SQL
0-shot Prompt 62.3 12.3 29.4 0.9
3-shot Prompt 44.3 10.9 35.7 1.3
SFT 82.1 63.2 38.0 0.4
Disambiguate-and-Parse
Interp. Prompt 81.8 26.0 81.9 16.9
w. Self-Correction 77.4 13.9 65.7 5.9
Gold Interp. SFT 87.4 61.2 62.6 0.3
Ours 92.3 53.2 84.4 18.8
Table 1: Single and Full Interpretation Coverage (%) on AmbiQT and Ambrosia datasets.
@@
As can be seen in Table 1, prompting (0-shot, 3-shot) performs poorly on ambiguous questions, which is consistent with the findings of Saparina and Lapata (2024a). Fine-tuning achieves excellent coverage on in-domain evaluation (AmbiQT), but fails to generalize to Ambrosia, which suggests that the model overfits specific patterns in AmbiQT. Interpretation generation (via prompting) shows promising results on both datasets compared to end-to-end methods. Self-correction filters out some valid interpretations. Fine-tuning on interpretations behaves very similarly to fine-tuning on SQL queries, but provides higher single-interpretation coverage, which suggests it is more accurate in predicting at least one correct interpretation.
@@
Interpretation generation with infilling further improves single and full interpretation coverage on both datasets. While it does not achieve the highest full coverage on AmbiQA, it substantially improves over end-to-end prompting reaching 53%, while delivering the best single interpretation coverage of 92%. It effectively generalises to new domains and ambiguity types, showing the best results on Ambrosia among all methods. However, the full coverage on Ambrosia is still relatively low (at 19%). Ambrosia is a challenging benchmark on its own but it is also possible that some interpretations are missed due to annotation differences between the two datasets.
@@
Text-to-SQL parsing is hard even with gold interpretations!
@@ 5 Related Work
More recently, various approaches have been proposed to correct systematic biases in LLM outputs. For example, Ji et al. (2024) propose a model-agnostic module that learns correctional residuals between preferred and dispreferred outputs. Welleck et al. (2023) use a corrector model to iteratively review imperfect generations from a base model. Similarly, critique generators can be developed using reinforcement learning Wadhwa et al. (2024a) or through fine-grained feedback Wadhwa et al. (2024b). Such correction approaches are most effective when guided by external tools (Kamoi et al., 2024). We follow this paradigm using a specialized infilling model to correct systematic LLM biases towards certain interpretations and validate our output through SQL execution.
@@ Appendix A AmbiQT and Ambrosia Examples
We provide samples from the AmbiQT and Ambrosia evaluation sets. AmbiQT is publicly available under the MIT license and Ambrosia is under the CC BY 4.0 license.
AmbiQT contains four types of ambiguity: column and table ambiguities, join ambiguity, and ambiguity due to precomputed aggregates. Ambrosia covers three types of ambiguities: scope, attachment, and vagueness. Note that column and table ambiguities from AmbiQT correspond to vagueness in Ambrosia, although they differ in the number of gold queries provided (2 in AmbiQT versus 3 in Ambrosia).
@@ Appendix F Self-Correction Prompt
Appendix G Training on AmbiQT and Ambrosia*
@@
Method Single Full Single Full
Text-to-SQL FT 81.0 63.1 83.2 69.5
Gold Interp. SFT 86.2 60.4 75.6 37.4
Ours 92.5 54.0 87.0 33.6
Table 6: Single and Full Interpretation Coverage (%), on AmbiQT and Ambrosia* test sets. Models are fine-tuned on both AmbiQT and Ambrosia* train. Symbol * denotes our split of the Ambrosia test set into training, development, and test.
Table 6 provides results when the models are trained on the maximum available data: the AmbiQT train set and re-split Ambrosia* training set. As both test sets are in-domain, we would not expect to see any advantages from our method in terms of full coverage. Nevertheless, it still provides the best single-interpretation coverage. Overall, the results in Table 6 are very similar to those obtained previously in the in-domain setting.

### v2

# Disambiguate First, Parse Later: Generating Interpretations for Ambiguity Resolution in Semantic Parsing

@@ Abstract
Handling ambiguity and underspecification is an important challenge in natural language interfaces, particularly for tasks like text-to-SQL semantic parsing. We propose a modular approach [^1] Our code and data can be downloaded from github.com/saparina/disambiguate-then-parse. that resolves ambiguity using natural language interpretations before mapping these to logical forms (e.g., SQL queries). Although LLMs excel at parsing unambiguous utterances, they show strong biases for ambiguous ones, typically predicting only preferred interpretations. We constructively exploit this bias to generate an initial set of preferred disambiguations and then apply a specialized infilling model to identify and generate missing interpretations. To train the infilling model, we introduce an annotation method that uses SQL execution to validate different meanings. Our approach improves interpretation coverage and generalizes across datasets with different annotation styles, database structures, and ambiguity types.
Disambiguate First, Parse Later: Generating Interpretations for
@@
Irina Saparina  and Mirella Lapata Institute for Language, Cognition and Computation, School of Informatics, University of Edinburgh 10 Crichton Street, Edinburgh EH8 9AB i.saparina@sms.ed.ac.uk   mlap@inf.ed.ac.uk
@@ 1 Introduction
While state-of-the-art large language models (LLMs) demonstrate remarkable performance on tasks like question answering, text-to-SQL parsing, and natural language inference, recent studies (Liu et al., 2023; Floratou et al., 2024) have shown they are not adept at handling ambiguity. They display systematic biases in their choice of interpretation (Kamath et al., 2024; Stengel-Eskin et al., 2024; Saparina and Lapata, 2024a), typically defaulting to a single interpretation when multiple ones exist.
@@
Figure 2: Disambiguate first parse later: we disambiguate database questions into natural language interpretations (by generating an initial set and filling in missing ones), and then a parser translates each interpretation into SQL.
@@ 2.1 Problem Formulation
In our text-to-SQL task, grammar $G$ defines valid SQL queries for a given database schema $S$. The database schema and table descriptions provide context $\mathcal{C}$ that can help disambiguate some queries. We focus on questions that remain ambiguous even when $\mathcal{C}$ is known, i.e., there exist multiple valid SQL queries $\{e_{1},...,e_{n}\}$ that respect $S$ and produce different result sets $\llbracket e_{i}\rrbracket$ when executed on the database. In the example in Figure 1, the question $u$ = “return the rating of each hotel” remains ambiguous for a given schema $S$ (see top table), as “rating” could refer to star ratings and guest reviews even when the database content is known.
@@ 2.2 Natural Language Interpretations for Explicit Disambiguation
We propose to resolve ambiguity prior to generating SQL expressions with natural language interpretations. More formally, for ambiguous utterance $u$, we first produce a set of unambiguous natural language interpretations $\{\hat{u}_{1},...,\hat{u}_{n}\}$, where each $\hat{u}_{i}$ captures one meaning of $u$. We then map each $\hat{u}_{i}$ to formal expression $e_{i}$. For example, given $u$ = “return the rating of each hotel”, our goal is to produce $\hat{u}_{1}$ = “How many stars were assigned to each hotel?”, $\hat{u}_{2}$ = “How did the customers review each hotel?” and $\hat{u}_{3}$ = “Show me the guest scores and star rating of each hotel” and map them to corresponding SQL queries $e_{1}$, $e_{2}$ and $e_{3}$ (see Figure 1). This approach has several advantages:
  • Unambiguous interpretations $\hat{u}_{i}$ are easier to translate into formal expressions, as they can be handled by standard semantic parsers (e.g., existing text-to-SQL models);
@@ 3.2 Evaluation Metrics
Our primary metric is Full Interpretation Coverage, which measures the proportion of examples where all valid SQL interpretations are present in the output. This metric aligns with the “All Found” metric from Ambrosia and “BothInTopK” from AmbiQT. Full Coverage is the most informative and relevant metric for our task, as it directly measures the system’s ability to recover the full set of available interpretations. We also report Single Interpretation Coverage, i.e., the proportion of examples where at least one valid interpretation is found (“EitherInTopK” from AmbiQT), as well as Precision and Recall. We compare SQL queries based on their execution results.
AmbiQT Ambrosia
End-to-End Text-to-SQL Single Full Recall Precision Single Full Recall Precision
0-shot Prompt 62.3 12.3 37.3 58.1 29.4 0.9 15.0 21.9
3-shot Prompt 44.3 10.9 27.6 33.0 35.7 1.3 17.5 21.3
SFT 82.1 63.2 72.7 74.3 38.0 0.4 20.0 29.4
AmbiQT Ambrosia
Disambiguate-and-Parse Single Full Recall Precision Single Full Recall Precision
Interp. Prompt 81.8 26.0 53.4 31.8 81.9 16.9 49.0 29.1
w. Self-Correction 77.4 13.9 45.7 46.0 65.7 5.9 34.5 29.4
Gold Interp. SFT 87.4 61.2 74.1 79.6 62.6 0.3 32.1 49.5
Ours 92.3 53.2 72.8 38.2 84.4 18.8 51.9 24.2
Table 1: Single and Full Interpretation Coverage, Recall and Precision (%) on AmbiQT and Ambrosia datasets.
@@ 4 Experimental Results
Table 1 is split into two sections. The first one presents end-to-end approaches, which attempt to directly predict multiple SQL queries for ambiguous questions. We report results for zero-shot prompting, few-shot prompting, and end-to-end fine-tuning (SFT). For prompting, we follow Saparina and Lapata (2024a) and explicitly instruct the model to generate multiple SQL queries if the question is ambiguous. For few-shot prompting, we sample 3 random examples from the corresponding dataset. The second section in Table 1 lists methods which use natural language interpretations to disambiguate first and then rely on text-to-SQL parsing of unambiguous questions. We report results for generating all possible interpretations through LLM prompting (which corresponds to our method without infilling), applying self-correction to this approach (details are in Appendix F), a fine-tuning method which is trained to generate reference interpretations instead of SQL queries, and our proposal which uses infilling to augment the set of default interpretations. We use an instruction-tuned Qwen2.5-Coder 32B for text-to-SQL generation.
@@
As can be seen in Table 1, prompting (0-shot, 3-shot) performs poorly on ambiguous questions, which is consistent with the findings of Saparina and Lapata (2024a). Fine-tuning achieves excellent results on in-domain evaluation (AmbiQT), but fails to generalize to Ambrosia, which suggests that the model overfits specific patterns in AmbiQT. Interpretation generation (via prompting) shows promising results on both datasets compared to end-to-end methods. Self-correction increases precision but also eliminates some valid interpretations, suggesting that the filtering task is non-trivial and requires further research. Fine-tuning on interpretations behaves very similarly to fine-tuning on SQL queries, but provides higher precision and single-interpretation coverage, which suggests it is more accurate in predicting at least one correct interpretation.
@@
Interpretation generation with infilling further improves single and full interpretation coverage on both datasets. While it does not achieve the highest full coverage on AmbiQA, it substantially improves over end-to-end prompting reaching 53%, while delivering the best single interpretation coverage of 92%. It effectively generalises to new domains and ambiguity types, showing the best results on Ambrosia among all methods across three out of four evaluation metrics. However, the full coverage on Ambrosia is still relatively low (at 19%). Ambrosia is a challenging benchmark on its own but it is also possible that some interpretations are missed due to annotation differences between the two datasets.
The results in Table 1 show that recall closely mirrors full coverage, offering little additional insight. Precision, on the other hand, can be misleadingly inflated when systems rarely predict all valid interpretations at once. For this reason, in subsequent experiments, we only report full and single interpretation coverage as the most informative metrics. See Appendix G for additional results.
@@
Error Type AmbiQT Ambrosia
Redundant interpr. 23.3 20.0
Overly general interpr. 16.7 20.0
Interpr. not matching intent 26.7 30.0
SQL generation errors 33.3 30.0
Table 4: Error types (%) on AmbiQT and Ambrosia.
Text-to-SQL parsing is hard even with gold interpretations!
Generalisation to new ambiguity types remains an open challenge.
To better understand our system’s limitations, we performed a manual error analysis on 50 randomly sampled examples with errors (30 from AmbiQT and 20 from Ambrosia). Table 4 summarizes the distribution of these errors. Redundant interpretations refer to near-paraphrases that differ slightly in structure. If their SQL queries yield the same result, we typically filter them out; however, minor variations (e.g., extra columns) can lead to execution mismatches, which are considered errors. Additionally, some interpretations remain overly vague, failing to fully disambiguate the input. These two error categories mostly originate from default interpretations ( 75%). We also observed SQL errors for valid interpretations, with 57% stemming from infilled interpretations, including incorrect joins and column mismatches. Syntax errors occurred in up to  20% of cases.
Most AmbiQT failures are due to join and aggregate ambiguities, which are not represented in the training data. Ambrosia presents even greater challenges with scope and attachment ambiguities, which remain particularly difficult.
@@ 5 Related Work
More recently, various approaches have been proposed to correct systematic biases in LLM outputs. For example, Ji et al. (2024) propose a model-agnostic module that learns correctional residuals between preferred and dispreferred outputs. Welleck et al. (2023) use a corrector model to iteratively review imperfect generations from a base model. Similarly, critique generators can be developed using reinforcement learning Wadhwa et al. (2024a) or through fine-grained feedback Wadhwa et al. (2024b). Such correction approaches are most effective when guided by external tools Kamoi et al. (2024). We follow this paradigm using a specialized infilling model to correct systematic LLM biases towards certain interpretations and validate our output through SQL execution.
@@ 7 Limitations
8 Acknowledgements
We thank the meta-reviewer and anonymous reviewers for their constructive feedback. We gratefully acknowledge the support of the UK Engineering and Physical Sciences Research Council (grant EP/W002876/1).
@@ Appendix A AmbiQT and Ambrosia Examples
We provide samples from the AmbiQT Bhaskar et al. (2023) and Ambrosia Saparina and Lapata (2024a) evaluation sets.
@@
AmbiQT contains column and table ambiguities, join ambiguity, and ambiguity due to precomputed aggregates. Ambrosia covers scope and attachment ambiguities, and vagueness. Note that column and table ambiguities from AmbiQT correspond to vagueness in Ambrosia, although they differ in the number of gold queries provided (2 in AmbiQT versus 3 in Ambrosia).
AmbiQT is publicly available under the MIT license and Ambrosia is under the CC BY 4.0 license.
@@ Appendix F Self-Correction Prompt
Appendix G Additional Results
Comparison with a larger-sized model
Table 7 extends Table 1 by including a larger baseline: the instruction-tuned Llama 3.1 70B model. While the 70B variant outperforms the 8B version overall, our system (based on the 8B model) still achieves higher coverage and recall. We expect that larger models could further improve results, however, this comes at a significantly higher computational cost.
Training on Ambrosia*
Table 8 extends Table 6 with recall and precision. Consistent with earlier findings, our approach achieves the highest coverage and recall on the AmbiQT test set, demonstrating strong out-of-domain generalization.
AmbiQT Ambrosia
Llama 3.1 70B Single Full Recall Precision Single Full Recall Precision
0-shot Prompt 76.6 26.5 51.5 65.8 32.3 0.0 14.1 32.0
3-shot Prompt 75.7 28.0 51.9 69.2 57.5 3.9 28.1 50.9
AmbiQT Ambrosia
Llama 3.1 8B Single Full Recall Precision Single Full Recall Precision
0-shot Prompt 62.3 12.3 37.3 58.1 29.4 0.9 15.0 21.9
3-shot Prompt 44.3 10.9 27.6 33.0 35.7 1.3 17.5 21.3
Ours 92.3 53.2 72.8 38.2 84.4 18.8 51.9 24.2
Table 7: Comparison of Llama 3.1 70B, Llama 3.1 8B and our approach on AmbiQT and Ambrosia datasets. Our approach is based on Llama 3.1 8B and fine-tuned on AmbiQT train.
@@ Appendix G Additional Results
Method Single Full Recall Precision Single Full Recall Precision
Text-to-SQL FT 56.1 10.2 33.2 52.7 77.1 66.4 71.6 73.3
Interp. Prompt 81.8 26.0 53.4 31.8 81.9 16.9 49.0 29.1
Gold Interp. SFT 81.9 4.1 43.0 79.5 71.0 38.9 53.7 61.5
Ours 88.0 30.0 58.5 35.3 87.8 30.5 59.7 26.3
Table 8: Single and Full Interpretation Coverage, Recall and Precision (%) on AmbiQT and Ambrosia* test sets. Models are fine-tuned on Ambrosia* train. Symbol * denotes our split of the Ambrosia test set into training, development, and testing.
AmbiQT Ambrosia*
Method Single Full Recall Precision Single Full Recall Precision
Text-to-SQL FT 81.0 63.1 72.1 73.3 83.2 69.5 76.2 78.1
Gold Interp. SFT 86.2 60.4 73.3 79.8 75.6 37.4 55.0 65.1
Ours 92.5 54.0 73.2 38.8 84.7 29.8 57.1 28.8
Table 9: Single and Full Interpretation Coverage, Recall and Precision (%) on AmbiQT and Ambrosia* test sets. All models are fine-tuned on both AmbiQT and Ambrosia* train. Symbol * denotes our split of the Ambrosia test set into training, development, and test.
Training on AmbiQT and Ambrosia*
Table 9 provides results when the models are trained on the maximum available data: the AmbiQT train set and re-split Ambrosia* training set. As both test sets are in-domain, we would not expect to see any advantages from our method in terms of full coverage. Nevertheless, it still provides the best single-interpretation coverage. Overall, the results in Table 9 are very similar to those obtained in the in-domain setting.